40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 2311     Entry time: Mon Nov 23 00:46:09 2009
Author: rana, rob 
Type: Update 
Category: PSL 
Subject: ISS RIN: Its too high by 10x 

This plot shows the RIN as measured by the ISS. Its ~2 x 10^-7, whereas its supposed to be more like 3 x 10^-8.

The ISS has DC coupled RIN channels (with a _F suffix) and AC coupled RIN channels (with a _FW suffix). By using a swept sine, Rob determined that the AC coupled channels have an AC coupling pole at ~80 Hz. The attached plot uses this and then has the overall gain adjusted to match with the _F channels below 10 Hz.

The _F channels can be converted directly into RIN by just dividing the spectra by the mean value of the time series. The dark offset of these channels is small and so this only introduces a ~5-10% calibration error.

Question #1: Why is the RIN so bad? According to the MEDM screen, the photocurrent on the MON/SENS PDs is 1.9/1.3 mA. That's sort of low, but should still allow us to get 5x10^-8 in RIN.

Question #2: Does it make an effect on the current DC Readout work? IF so, should we try to fix up the ISS in a temporary way? Since the in-loop and out-of-loop detectors are completely coherent, all of the noise is likely just unsuppressed noise from the laser. We are unable to increase the gain because of the high frequency noise from the NPRO.

 

Let's remember to replace this ISS with a new one that can drive an AOM. Need a volunteer to get us a new ISS.

 

Attachment 1: Untitled.png  44 kB  | Hide | Hide all
Untitled.png
ELOG V3.1.3-